OpenStudy (anonymous):

how to solve for :

OpenStudy (anonymous):

\[\log_{4.3301-2.5i} (2+3i)\]

OpenStudy (alekos):

this is a tricky one!

ganeshie8 (ganeshie8):

Easy - change into exponential form and the log cancels out!

ganeshie8 (ganeshie8):

\[\log_{4.3301-2.5i} (2+3i) = \dfrac{\ln (2+3i)}{\ln(4.3301 - 2.5i)} = \dfrac{\ln (re^{??})}{\ln(se^{??})}\]

ganeshie8 (ganeshie8):

wolfram is your best friend in verifying your answer : http://www.wolframalpha.com/input/?i=log_%284.3301-2.5i%29+%282%2B3i%29

OpenStudy (anonymous):

thank you! i don't know how to use that thing ill try harder!!! :DD

ganeshie8 (ganeshie8):

wolfram gives you the general answer but looks like you're working with principal angles so you will still need to work few things manually i guess :)

OpenStudy (alekos):

very interesting. I get a real answer?

ganeshie8 (ganeshie8):

wolfram says 0.54 + 0.79i

ganeshie8 (ganeshie8):

this has some info about taking log of a complex number : http://en.wikipedia.org/wiki/Complex_logarithm idk what it means in real though :P

OpenStudy (alekos):

I've just done it by hand and you're right. fascinating problem, thank you.

OpenStudy (alekos):

karla. do you want to see the real maths?

OpenStudy (anonymous):

:) yes